Gustav’s father is seven times as old as he is. One year ago, he was nine times as old as Gustav. Find Gustav’s present age.

Answer by richwmiller(17219)

f=7g
f-1=9*(g-1)
7g-1=9g-9
8=2g
g=4
f=28
check
27=9*3
ok